# Veli: Unsupervised Method and Unified Benchmark for Low-Cost Air Quality Sensor Correction

### v1

@@ 4.3 Correction Results
Figure 5 presents a 12-hour-average time series over a two-month period in Utrecht, which has the worst raw LCS accuracy among the selected regions. The region in the graph between the red-dashed lines highlights a four-day window, whose hourly sampling was shown earlier in Figure 1. Veli successfully captures both short- and long-term trends and spikes, despite our model being completely reference-free. While our model performs best when the raw LCS data exhibits an underlying trend (i.e., is not completely random), it is also designed to handle a common failure mode in which sensors produce random readings. We demonstrate this robustness in Section 4.4.
@@ 4.4 Model Analysis and Discussion
PM2.5 readings typically exhibit strong autocorrelation that gradually decays over time, primarily driven by the underlying pollutant concentrations (Zaini et al. 2022). In addition to this inherent structure, noise from LCS can also introduce significant autocorrelation, often persisting for up to 48 hours, as illustrated in Figure 7. The raw LCS readings remain highly autocorrelated for an extended period of time, in contrast to the trends seen in the reference stations. Our model eliminates this noise, producing outputs that closely match the reference time series. This behavior is consistent with the discussion in Section 3.1, showing that our model corrects the readings without compromising temporal information.
@@
To simulate adversarial sensor failure (channel dropout), we took the original test data and randomly replaced a fixed number of the 10 sensor readings, $n$, with ‘NA’ values for each hourly sample. We then evaluated the model’s performance for different values of $n$, from 1 to 9. Figure 8 shows how these injected failures degrade the correction’s performance, but remains within an acceptable range of accuracy (MAE $<$ 10). For sensors that are already accurate (e.g., Nijmegen’s LCS), using a variation of Veli with fewer channels would be beneficial, which we show in the next subsection and further in Appendix B.1.
@@ A.1 Standard ELBO Identity
Following the standard VAE formulation (Kingma and Welling 2014), we take the logarithm and introduce the variational posterior to define the ELBO. The objective is to maximize this bound to obtain a tractable lower bound on the log-likelihood of the data, such that:
@@ A.2 KL penalties
=-= - $\mathbb{E}_{q_{\phi}(z|x)}\left[D_{\rm KL}(q_{\theta}(y|z,x)\|p(y|z))\right]$
@@
[…]
@@ A.3 ELBO with Reconstruction Term
Following (Higgins et al. 2017), we introduced three hyperparameters $\alpha,\beta_{\rm z}$, and $\beta_{\rm y}$ to weight each term. The final minimization target, the negative ELBO, becomes:
@@ B.2 Predictions with Credible Intervals
A noticeable limitation of our proposed method is underestimation of the true reference values. Figure 13 shows the distributions of raw LCS readings, Veli’s outputs, and reference values. Although our model successfully closely approximates the underlying readings, there is a pattern of underestimation when compared to the reference readings. This underestimation is a consequence of the unsupervised approach. Unlike previous reference-based works that use reference readings to correct LCS readings (Li et al. 2021), Veli does not have a ground-truth signal to guide the output distribution toward the true data distribution.
@@ C.1 Dataset Building
AQ-SDR aggregates the LCS data from three major citizen-science initiatives: SamenMeten, Sensor.Community, and Location Aware Sensing System (LASS) community (Chen et al. 2017). SamenMeten (Measure Together) is a Dutch citizen-science platform that supports public participation in environmental monitoring, focusing on air, water, and noise quality. Sensor.Community is a global citizen-science network that generates open AQ and noise data via LCS. Lastly, LASS is a Taiwanese citizen-science platform that leverages citizen-contributed AQ data and environmental sensing networks.
@@ C.2 Preprocessing
The dataset and the code to identically replicate the preprocessing will be made publicly available. First, we resample every sensor’s data using its hourly average, since different devices provide readings at different temporal frequencies. We then define a range-validation step that applies hard bounds for all values for each pollutant and meteorological measurement. For example, we drop outdoor temperature readings outside the range [-50C, 70C], as they are implausible in an inhabited area. Next, we split every data stream into batches of two months and pass it through a density-based outlier detection (DBSCAN) with a lenient threshold (Ester et al. 1996). The purpose of this outlier detection is to eliminate long periods of abnormal readings. Outliers indicating potential sensor failure are set to ‘NA’ (e.g., an extended period of PM2.5 values near 600 $\mu\mathrm{g}/\mathrm{m}^{3}$). After this preprocessing step, we ensure that every sensor was functional (non-NA hours) for at least 6,000 hours (65% of a full year), though not necessarily consecutive. This step prevents data streams from being empty or overly sparse. The code and further details on preprocessing are documented and will be shared along with the dataset.

### v2

@@ 4.3 Correction Results
Figure 5 presents a 12-hour-average time series over a two-month period in Utrecht, which has the worst raw LCS accuracy among the selected regions. The region in the graph between the red-dashed lines highlights a four-day window, whose hourly sampling was shown earlier in Figure 1. Veli successfully captures both short- and long-term trends and spikes, despite our model being completely reference-free. While our model performs best when the raw LCS data exhibits an underlying trend (i.e., is not completely random), it is also designed to handle a common failure mode with random readings, demonstrated in Section 4.4.
@@ 4.4 Model Analysis and Discussion
PM2.5 readings typically exhibit strong autocorrelation that gradually decays over time, primarily driven by the underlying pollutant concentrations (zaini_pm25_2022). In addition to this inherent structure, noise from LCS can also introduce significant autocorrelation, often persisting for up to 48 hours, as illustrated in Figure 7. The raw LCS readings remain highly autocorrelated for an extended period of time, in contrast to the trends seen in the reference stations. Our model eliminates this noise, producing outputs that closely match the reference time series. This behavior is consistent with the discussion in Section 3.1, showing that our model corrects the readings without compromising temporal information.
@@
To simulate adversarial sensor failure (channel dropout), we took the original test data and randomly replaced a fixed number of the 10 sensor readings, $n$, with ‘NA’ values for each hourly sample. We then evaluated the model’s performance for different values of $n$, from 1 to 9. Figure 8 shows how these injected failures degrade the correction’s performance, but remains within an acceptable range of accuracy (MAE << 10). For sensors that are already accurate (e.g., Nijmegen’s LCS), using a variation of Veli with fewer channels would be beneficial, which we show in the next subsection and further in Appendix B.1.
@@ 5 Conclusion
6 Acknowledgements
We sincerely thank the Dutch government for supporting this research with the starter grant (startersbeurzen). We also thank the organizations and researchers who provide the open data to enable this research, including the Dutch National Institute for Public Health and the Environment (RIVM), the Dutch Royal Netherlands Meteorological Institute (KNMI), Dr. Ling-Jyh Chen in Taiwan Academia Sinica for the AirBox project, the Taiwan Ministry of Environment, the Sensor.Community platform, and the European Environmental Agency (EEA). We also thank the GGD Amsterdam and RIVM for providing information about how air quality sensor stations work in the Netherlands. We also thank the CREATE Lab at the Robotics Institute at Carnegie Mellon University for the technical support in building the air quality dashboard.
@@ A.1 Standard ELBO Identity
Following the standard VAE formulation (vae), we take the logarithm and introduce the variational posterior to define the ELBO. The objective is to maximize this bound to obtain a tractable lower bound on the log-likelihood of the data, such that:
@@ A.2 KL penalties
$=$- $\mathbb{E}_{q_{\phi}(z|x)}\left[D_{\rm KL}(q_{\theta}(y|z,x)\|p(y|z))\right]$
@@
In our implementation, we sample latent variables $z$ and $y$ using the reparameterization trick similar to VAEs. We also sample multiple values of $(z,y)$ such that the expectations $\mathbb{E}_{\rm q_{\rm\theta,\phi}(y,z\mid x)}$ and $\mathbb{E}_{\rm q_{\phi}(z\mid x)}$ are estimated per batch, allowing us to omit the expectation in the final ELBO expression. Combining these terms, we obtain the final ELBO term:
@@ A.3 ELBO with Reconstruction Term
Following (betavae), we introduced three hyperparameters $\alpha,\beta_{\rm z}$, and $\beta_{\rm y}$ to weight each term. The final minimization target, the negative ELBO, becomes:
@@ B.2 Predictions with Credible Intervals
A noticeable limitation of our proposed method is underestimation of the true reference values. Figure 13 shows the distributions of raw LCS readings, Veli’s outputs, and reference values. Although our model successfully closely approximates the underlying readings, there is a pattern of underestimation when compared to the reference readings. This underestimation of approximately 3 $\mu\mathrm{g}/\mathrm{m}^{3}$ is observed only for air quality readings below 15 $\mu\mathrm{g}/\mathrm{m}^{3}$, an error that is not significant for hazardous event prediction. We also point out that our proposed credible intervals provide a robust mitigation strategy against this limitation. This underestimation is a consequence of the unsupervised approach. Unlike previous reference-based works that use reference readings to correct LCS readings (maximizingcorrelation), Veli does not have a ground-truth signal to guide the output distribution toward the true data distribution.
@@ C.1 Dataset Building
AQ-SDR aggregates the LCS data from three major citizen-science initiatives: SamenMeten, Sensor.Community, and Location Aware Sensing System (LASS) community (taiwan_data). SamenMeten (Measure Together) is a Dutch citizen-science platform that supports public participation in environmental monitoring, focusing on air, water, and noise quality. Sensor.Community is a global citizen-science network that generates open AQ and noise data via LCS. Lastly, LASS is a Taiwanese citizen-science platform that leverages citizen-contributed AQ data and environmental sensing networks.
@@ C.2 Preprocessing
The dataset and the code to identically replicate the preprocessing will be made publicly available. First, we resample every sensor’s data using its hourly average, since different devices provide readings at different temporal frequencies. We then define a range-validation step that applies hard bounds for all values for each pollutant and meteorological measurement. For example, we drop outdoor temperature readings outside the range [-50C, 70C], as they are implausible in an inhabited area. Next, we split every data stream into batches of two months and pass it through a density-based outlier detection (DBSCAN) with a lenient threshold (dbscan). The purpose of this outlier detection is to eliminate long periods of abnormal readings. Outliers indicating potential sensor failure are set to ‘NA’ (e.g., an extended period of PM2.5 values near 600 $\mu\mathrm{g}/\mathrm{m}^{3}$). After this preprocessing step, we ensure that every sensor was functional (non-NA hours) for at least 6,000 hours (65% of a full year), though not necessarily consecutive. This step prevents data streams from being empty or overly sparse. The code and further details on preprocessing are documented and will be shared along with the dataset.